# The PLATO Mission

### v1

@@ 8.1 Generation of the Data Products
The scientific specifications for the on-ground software for generating the data products are provided by the PSM, which benefits from extensive expertise from the European scientific community. These specifications account for the lessons learned from CoRoT, Kepler, K2, CHEOPS, and TESS and build on the latest advances in stellar and exoplanet sciences. During operations, the PSM will scientifically review the data processing chain and provide updated scientific specifications for algorithms and tools where needed. For each target field, PSM provides the requirements for the PIC. PSM also includes the Ground-Based Observing Programme (GOP), which will provide the Lg Data under PSM responsibility, including the observations needed for the confirmation of the planets and the radial velocities for the determination of their masses. The PSM will also scientifically validate the final list of planets and their characteristics. The PCOT will support the calibration activities and operations, e.g. by providing input to the procedures needed for payload operation and for scientific mission planning.
The PDC responsibilities include the definition and implementation of the L1 and Calibration (CPDS) pipeline (to be run by SOC at ESAC, Spain) and the development of the software for processing the L2 and L3 data products (L3 combines L2 and Lg). The generation of the PLATO data products will be performed within two main pipelines, the Exoplanet Analysis Pipeline (EAS) and the Stellar Analysis Pipelines (SAS). These pipelines do not work independently, and several intermediate products (e.g. stellar rotation periods; transit detections) will be interchanged between them. A more detailed description of these pipelines will be provided elsewhere after finalising their design.
@@ 8.2 Data Products
Level-0: Depending on target, Level-0 data include unprocessed imagettes or on-board pre-processed light curves and centroid curves. Level-0 data are downloaded for the respective targets from all cameras.
Level-1: Calibrated light curves and centroid curves for all targets and cameras produced on-ground. Level-1 data include, e.g., corrections for instrumental effects and the light curves and centroid curves derived from imagettes. In addition, for the normal cameras, camera-averaged Level-1 light curves and centroid curves are provided for each target star. For the fast cameras, no average light curves are produced, so that the colour information is preserved.
@@ 8.3 Data Releases
The timeline for the releases of Level-0, Level-1, and Level-2 products has been specified with a cadence of three months. This is the time interval between two 90 rotations of the spacecraft, which are necessary to keep the adequate orientation with respect to the Sun. After each three months data acquisition period, three months will be required for the validation of the Level-1 products by the Science Operations Centre and by the PMC. Exceptionally, six months will be needed for Level-1 product validation of the first three months of nominal mission data. This longer validation period is justified because the data pipelines are expected to undergo the most important updates, resultant from the system calibration and characterisation with the first on-board data.
@@
For the 24,0000 non-prime sample targets, Level-0, Level-1 and Level-2 products for each observing quarter will be publicly released within three months of the corresponding Level-1 product validation period.
@@ 9 The Ground-based Observations Programme
The Ground-based Observation Programme (GOP) within the PMC will play an important role for the PLATO mission: it will be in charge of organising the ground-based observations needed to confirm the planetary nature of PLATO transit candidates. PLATO follows a strategy first established by the CoRoT mission by setting up a coordinated approach to ground-based follow-up managed by the mission consortium in order to secure sufficient resources and efficient planning of ground-based activities.
The PMC will establish and manage the GOP team with contributions from international partners taking part in the follow-up. The GOP Team will respond to calls issued by ground-based facilities (e.g., ESO) to grant observing time for the confirmation and characterisation of the candidates in the “prime sample” (see Section 6). Dedicated facilities taking part, e.g., in the filtering process of planet candidates will also contribute to PLATO. A key planet parameter for the confirmation and characterisation of planets is their mass, which can be determined by means of high-resolution spectroscopic observations providing precise radial velocities. But additional photometric and low-resolution facilities will as well be useful to discard false positive scenarios.
Within the GOP team tools are being created to interface to the database of PLATO candidates for the filtering of candidates to detect astrophysical mimics and the subsequent radial velocity follow-up of the planet candidates. For each object, observations will be optimised to exploit the available resources. Facilities taking part in the GOP will also have to demonstrate their performance accuracy and hence the observations that they can be most efficiently deployed to carry out. The GOP also has significant data quality control measures to ensure that only data of sufficient quality is used in the decision making process and eventually in the system modelling.
@@
The GOP is also responsible for gathering any preparatory observations being attempted (e.g. detailed spectroscopic observations of the PLATO stellar samples). For target samples beyond the propriety and prime samples, the GOP will perform and support preparatory and follow-up observations on a best effort basis and provide, together with the PDC, an interface to data obtained on PLATO targets by the general community.
@@ 10 The Guest Observer Programme
The PLATO Guest Observer (GO) programme will allow all interested groups and individuals to submit proposals to observe targets which are not included in the PLATO prime sample (see Section 8.2) and have science goals which are not covered by the PLATO core science objectives. Examples of complementary science cases which could be addressed via the GO programme are discussed in Section 5. An average of 8% of the science data rate (excluding calibration data), averaged over the mission lifetime, will be allocated to the GO programme. This will allow for an extended complementary science programme, while preserving the resources for an optimal observation of the core science targets. The total number of GO objects that may be observed with this allocation will range from thousands to tens of thousands, depending on the sampling times and whether imagettes or light curves are requested. ESA will appoint an independent Time Allocation Committee (TAC) for the evaluation and selection of the proposals based on scientific merit.
@@
In addition to the science done by the successful GO applicants, a general variability classification will be performed on all PLATO data once they are released publicly. This probabilistic classification will be based on applications of supervised and unsupervised machine learning methods, resulting in an electronic public PLATO variability catalogue, which will be updated regularly as new data will be released to the worldwide community.
@@ 11 Expected Instrument Performance
The design of the PLATO mission is driven by the need to obtain a core sample which is dominated by stellar flux white noise and where instrumental effects play only a minor role. Although the nominal mission operation time is 4.25 years at this point, the payload is designed to maintain a level of performance that allows the science objectives to be achieved up to an extended mission end-of-life (EOL) of 6.5 years. The EOL scenario assumes that up to two cameras could be lost during the mission due to technical failures. This number results from a reliability analysis performed during the design phase.
@@
At the time of writing this paper, instrument requirements for PLATO have been settled and a number of simulations of the expected instrument performance have been performed. Figure 21 illustrates the importance of different instrumental noise sources for PLATO (see Börner et al. in prep.). Throughout the main target magnitude range of PLATO (8 mag to 11 mag) photon noise dominates the signal over instrumental effects. At the very bright end, jitter noise from the satellite dominates. Below about 12.5 mag electronic read-out noise starts to dominate.
Stars brighter than magnitude 8 will saturate with the N-CAMs (brighter than 4.5 with the F-CAMs). However, it shall be possible to extract precise photometry from saturated targets. For slightly saturated stars (around magnitude 8) no extension of the nominal window of 6x6 pixels will be necessary. For moderately saturated targets (between magnitudes 4 and 8) saturated star descriptors are needed with the N-CAMs while smearing photometry will be possible with the F-CAMs. Finally, for the few highly saturated targets (expected 10 per camera) smearing photometry is technically possible, but alternatives are being considered [e.g. 317].
The noise-to-signal simulator PINE (Börner et al. in prep) is using the current best knowledge of the instrument design. PINE values were used, e.g., for the development of the PIC. The detailed simulation of stellar noise-to-signal ratios (NSR)s across the PLATO FoV is shown in Fig. 21 versus stellar magnitude. As can be seen, the required P1 noise performance of 50 ppm in 1 hour is met in the center of the F0V (24 cameras) for stars brighter than 11 mag. At the corners (6 cameras) the same performance is reached for stars brighter than 10 mag. Overall, when applied to real sky fields, the noise performance of PLATO is sufficient to provide the required number of stars in the respective P1 to P5 samples.
@@ 12.2 Launch and operations
PLATO is planned to be launched with an Ariane 62 rocket by the end of 2026. Since the PLATO payload’s line of sight is up-oriented towards the zenith, and blinding and illumination of the cameras by the Sun must be avoided, the strategy selected for launch and transfer includes an intermediate circular parking Low-Earth orbit. In addition, the upper stage of Ariane 62 provides a barbeque mode, such that, it can rotate around its longitudinal axis during its coasting phase, avoiding illumination of the payload. By adopting this strategy, launch is possible all year round with some exclusion windows to avoid Moon and Earth eclipses during transfer. PLATO will be injected from the parking orbit into an eclipse-free Lissajous orbit around the Earth-Sun Lagrangian point 2 (L2), located 1.5 million km from Earth. During the transfer phase, which will take about 30 days, the commissioning phase will start and will run until the check-out and calibration of the spacecraft and its payload are completed, at the latest three months after launch. At that time the routine science operations will begin with a nominal duration of 4 years. Extension of the scientific operations will be possible as the satellite is being built and verified for an in-orbit lifetime of 6.5 years and will accommodate consumables for at least 8.5 years.
@@ 13.1 General payload description
The PLATO payload has two major science drivers constraining its design. Compared to previous missions like Kepler, PLATO has been optimised to observe stars bright enough to allow radial velocity measurements on ground to complement the photometric measurements in space. Since such bright stars are relatively scarce on the sky, a very large total field of view of 2132 deg2 was required. Furthermore, to be able to observe in a given target field the brightest as well as fainter stars, the instrument photon collecting area is split into 26 individual cameras. We call the 24 cameras which are used exclusively for science operation "normal" cameras and the 2 additional cameras, used also for fine-pointing on top of science operation, "fast" cameras. With this configuration PLATO will be able to obtain very stable, high precision, long-time photometric data of a large number of bright stars. Combining the information provided by the planetary transits, with the stellar data derived by asteroseismology techniques and the radial velocity measurements obtained from the ground, PLATO will provide a high-precision characterisation of the radius, mass and age of exoplanetary systems.
@@
Table 7 summarises the main data of the instrument. Each 20 cm class camera (12 cm entrance pupil) has a total field of view (FoV) of 1037 square degrees [see 248]. The 24 normal cameras are organised on the optical bench of the spacecraft in 4 groups of 6 co-aligned cameras each. The sky fields of the four camera groups are offset with respect to a central line of sight by 9.2 degrees, pointing towards the four corners of a square of 13.0 deg, in order to widen the total overlapping FoV of the instrument. With this approach the number of brightest stars within SNR requirements is augmented with respect to a full coalignment of the whole set of cameras. Figure 25 illustrates how many cameras observe the same part of the target field for a pointing. The total instrument field therefore is about 2132 deg² (with rounded edges). Figure 26 shows the Structural and Thermal Model (STM) of the payload during integration at OHB, ready for the qualification activities of the Payload Module.
@@
| Total Field of View | 2132 deg² |
@@
| Attitude Pointing Drift Error (PDE) | 3 arcsec half cone angle over 3 months in the directional direction |
| 6 arcsec half cone angle over 3 months in the rotational direction | |
| Attitude Pointing Repeatability Error (PRE) | 3 arcsec half cone angle in the directional direction |
| 6 arcsec half cone angle in the rotational direction | |
| Relative Attitude Pointing Error (RPE) | 1 arcsec half cone angle over 25s in the directional direction |
| 2 arcsec half cone angle over 25s in the rotational direction | |
@@
| Instrument peak power consumption | 890 W |
@@
| Normal electronic chain | 12 Data Processing Units (DPUs): 1 for 2 normal cameras |
| 2 normal Ancillary Electronic Units (AEU): 1 for 12 normal cameras | |

### v2

@@ 8.1 Generation of the Data Products
The scientific specifications for the on-ground software for generating the data products are provided by the PSM, which benefits from extensive expertise from the European scientific community. These specifications account for the lessons learned from CoRoT, Kepler, K2, CHEOPS, and TESS and build on the latest advances in stellar and exoplanet sciences. During operations, the PSM will scientifically review the data processing chain and provide updated scientific specifications for algorithms and tools where needed. For each target field, PSM provides the requirements for the PIC. PSM also includes the Ground-Based Observing Programme (GOP), which will provide the Lg Data (ground-based follow-up) under PSM responsibility, including the observations needed for the confirmation of the planets and the radial velocities for the determination of their masses. The PSM will also scientifically validate the final list of planets and their characteristics. The PCOT will support the calibration activities and operations, e.g. by providing input to the procedures needed for payload operation and for scientific mission planning.
The PDC responsibilities include the definition and implementation of the L1 and Calibration (CPDS) pipeline (to be run by SOC at ESAC, Spain) and the development of the software for processing the L2 and L3 data products (L3 combines L2 and Lg). See Section 8.2 for a detailed description of PLATO´s L0-L3 and Lg data products. The generation of the data products will be performed within two main pipelines, the Exoplanet Analysis Pipeline (EAS) and the Stellar Analysis Pipelines (SAS). These pipelines do not work independently, and several intermediate products (e.g. stellar rotation periods; transit detections) will be interchanged between them. A more detailed description of these pipelines will be provided elsewhere after finalising their design.
@@ 8.2 Data Products
Level-0: Depending on target, Level-0 data include unprocessed imagettes or on-board pre-processed light curves and centroid curves. Level-0 data are downloaded for the respective targets from all cameras. The concept for the optimal aperture used for the in-flight photometric extraction is given in [223].
Level-1: Calibrated light curves and centroid curves produced on-ground for all targets and cameras. Level-1 data include, e.g., corrections for instrumental effects and the light curves and centroid curves derived from imagettes. In addition, for the normal cameras, camera-averaged Level-1 light curves and centroid curves are provided for each target star. For the fast cameras, no average light curves are produced, so that the colour information is preserved.
@@ 8.3 Data Releases
The timeline for the releases of Level-0, Level-1, and Level-2 products has been specified with a cadence of three months. This is the time interval between two 90 rotations of the spacecraft, which are necessary to keep the adequate orientation with respect to the Sun. After each three months data acquisition period, three months will be required for the validation of the Level-1 products by the Science Operations Centre and by the PMC. Exceptionally, six months will be needed for Level-1 product validation of the first three months of nominal mission data. This longer validation period is justified because the data pipelines are expected to undergo the most important updates, resultant from the system calibration and characterisation with the first on-board data. Note also that reprocessing of the data will take place as needed due to algorithm improvements, and at the end of each pointing when the full data set is available.
@@
For the 240,000 non-prime sample targets, Level-0, Level-1 and Level-2 products for each observing quarter will be publicly released within three months of the corresponding Level-1 product validation period.
@@ 9 The Ground-based Observations Programme
The Ground-based Observation Programme (GOP) within the PMC plays an important role for the PLATO mission. It is in charge (for the "prime" and "proprity" samples) of organising and performing the ground-based observations needed to confirm the planetary nature of PLATO transit candidates and determine planet and stellar parameters at the required precision to reach the goals of the mission. The PMC will establish and manage the GOP team with contributions from international European partners taking part in the follow-up. The GOP Team will respond to calls issued by ground-based facilities (e.g., ESO) to grant observing time for the confirmation and characterisation of the candidates in the “prime sample” (see Sect. 6).
A key planet parameter for the confirmation and characterisation of planets is their mass, which can be determined by means of high-resolution spectroscopic observations providing precise radial velocities (RV) as e.g. the HARPS or Carmenes spectrographs. For terrestrial planets in the habitable zone of their stars, a precision at the level of a few 10s of cm/s is required that can only be achieved by the new generation of high-resolution super-stable spectrographs as e.g. ESPRESSO on the ESO VLT. In addition to RV measurements, photometric, imaging and lower-resolution spectroscopy facilities will be very useful as well to discard false positive scenarios. In this context, PLATO measurements of stellar rotation and activity cycles (see Sect. 4.6) will be useful to mitigate the impact of stellar activity on the radial velocity follow-up and eliminate false positive detection of non-transiting planets arising from activity.
The PLATO consortium is following a strategy first established by the CoRoT mission by setting up a coordinated approach to ground-based follow-up managed by the mission consortium in order to secure sufficient resources and efficient planning of ground-based activities. As mentioned above, this will involve extensive ground-based observations of different types and on a variety of facilities, including the largest telescopes with instruments capable of providing the most precise RV measurements. Dedicated smaller facilities taking part, e.g., in the filtering process of planet candidates will also contribute to PLATO. Observations have therefore to be scheduled in an optimal way, making a proper match between facilities and the observation needs, avoiding unnecessary duplications while ensuring a coherent use of the facilities registered in the GOP to obtain data at the required level of quality. Quality checks of the data will be performed to ensure that only data of sufficient quality is used in the decision making process and eventually in the system modelling, before possibly conducting additional observations when needed. Existing data in available archives will be used as well, and reprocessed when needed. The GOP data products (new or reprocessed) are delivered to the PDC where PLATO light curves are processed together with the ground-based data and where the stellar and planetary parameters are derived.
@@
The GOP is also responsible for gathering any preparatory observations being attempted (e.g. multiplex spectroscopic observations of stars in the PLATO field). For target samples beyond the propriety and prime samples, the GOP will perform and support preparatory and follow-up observations on a best effort basis taking advantage of momentarily available unused facilities and the GOP operational infrastructure developed for the mission. It is also foreseen to use data obtained on PLATO targets by the general community and generously provided to the PMC.
Although the GOP organisation and subsystems are designed to work in an automatic way, human intervention is still necessary in occasional cases, in order to have a smooth flow of information and data between participating facilities and the PDC, with a proper allocation of the needed resources. Teams are foreseen to be established to address GOP-related matters at the scientific and operational levels.
The participation to the GOP activities is fully open to the astronomers from ESA member states who will actively involve themselves in the project. They usually become full consortium members with the same rights and obligations as the other PMC members (strictly following the PLATO data access and publication policies). The participation of colleagues from outside of ESA is considered on a case by case basis. In order to allow for an optimal though flexible programming and use of facilities, it will also be possible to have actual observers at the telescopes who are not PLATO members but are active in time sharing coordination of projects run on specific facilities. Facilities taking part in the GOP are registered with the relevant associated information in a dedicated database maintained by the GOP. By facility we mean telescope+instrument (including an efficient data reduction software). The main a priori constraints they will have to fulfil are i) to be accessible by GOP members for extended periods of time, and ii) to have demonstrated their performance accuracy (through published results or benchmark tests), so they can be deployed for observations in the most efficient way.
@@ 10 The Guest Observer Programme
The PLATO Guest Observer (GO) programme will allow all interested groups and individuals to submit proposals to observe targets which are not included in the PLATO prime sample (see Section 8.2) and have science goals which are not covered by the PLATO core science objectives. A detailed description of the scope of the programme and of its boundaries with respect to the core science will be outlined in the GO programme policies and procedures that are currently in preparation. A few examples of complementary science cases which could be addressed via the GO programme are discussed in Section 5. An average of 8% of the science data rate (excluding calibration data), averaged over the mission lifetime, will be allocated to the GO programme. This will allow for an extended complementary science programme, while preserving the resources for an optimal observation of the core science targets. The total number of GO objects that may be observed with this allocation will range from thousands to tens of thousands, depending on the sampling times and whether imagettes or light curves are requested. ESA will appoint an independent Time Allocation Committee (TAC) for the evaluation and selection of the proposals based on scientific merit.
@@ 11 Expected Instrument Performance
The design of the PLATO mission is driven by the need to obtain a core sample which is dominated by stellar flux white noise and where instrumental effects play only a minor role. Although the nominal mission operation time is 4.25 years at this point, the payload is designed to maintain a level of performance that allows the science objectives to be achieved up to an extended mission end-of-life (EOL) of 6.5 years. The EOL scenario assumes that up to two cameras could be lost during the mission due to technical failures (i.e. 22 operating N-CAMs instead of the 24 N-CAMs available beginning-of-life, BOL). This number results from a reliability analysis performed during the design phase. Along the text, we are using the EOL as conservative worst-case scenario for mission performance, unless stated differently. For example, the targets of the PLATO Input Catalogue are chosen according to their BOL performance. This is the most informative choice for the scientific community and representative of the performance that will be found at the beginning of the mission. However, the planet yield and stellar characterization performances are computed with EOL performance, to make sure that the scientific goals of the project are reached in a realistic worst-case scenario.
@@
At the time of writing this paper, instrument requirements for PLATO have been settled and a number of simulations of the expected instrument performance have been performed. Figure 21 illustrates the importance of different instrumental noise sources for PLATO (see Börner et al [47]). Throughout the main target magnitude range of PLATO (8 mag to 11 mag) photon noise dominates the signal over instrumental effects. At the very bright end, jitter noise from the satellite dominates. Below about 12.5 mag electronic read-out noise starts to dominate.
Stars brighter than magnitude 8 will saturate with the N-CAMs (brighter than 4.5 with the F-CAMs). However, it shall be possible to extract precise photometry from saturated targets. For slightly saturated stars (around magnitude 8) no extension of the nominal window of 6x6 pixels will be necessary. For moderately saturated targets (between magnitudes 4 and 8) saturated star descriptors are needed with the N-CAMs while smearing photometry will be possible with the F-CAMs. Finally, for the few highly saturated targets (expected 10 per camera) smearing photometry is technically possible, but alternatives are being considered [e.g. 341].
The noise-to-signal simulator PINE ( Börner et al [47]) is using the current best knowledge of the instrument design. PINE values were used, e.g., for the development of the PIC. The detailed simulation of stellar noise-to-signal ratios (NSR)s across the PLATO FoV is shown in Fig. 21 versus stellar magnitude. As can be seen, the required P1 noise performance of 50 ppm in 1 hour is met in the center of the F0V (24 cameras) for stars brighter than 11 mag. At the corners (6 cameras) the same performance is reached for stars brighter than 10 mag. Overall, when applied to real sky fields, the noise performance of PLATO is sufficient to provide the required number of stars in the respective P1 to P5 samples.
@@ 12.2 Launch and operations
PLATO is planned to be launched with an Ariane 62 rocket by the end of 2026. Since the PLATO payload’s line of sight is up-oriented towards the zenith, and since blinding and illumination of the cameras by the Sun must be avoided, the strategy selected for launch and transfer includes an intermediate circular parking Low-Earth orbit. In addition, the upper stage of Ariane 62 provides a barbeque mode, such that, it can rotate around its longitudinal axis during its coasting phase, avoiding illumination of the payload. By adopting this strategy, launch is possible all year round with some exclusion windows to avoid Moon and Earth eclipses during transfer. PLATO will be injected from the parking orbit into an eclipse-free Lissajous orbit around the Earth-Sun Lagrangian point 2 (L2), located 1.5 million km from Earth. During the transfer phase, which will take about 30 days, the commissioning phase will start and will run until the check-out and calibration of the spacecraft and its payload are completed, at the latest three months after launch. At that time the routine science operations will begin with a nominal duration of 4 years. Extension of the scientific operations will be possible as the satellite is being built and verified for an in-orbit lifetime of 6.5 years and will accommodate consumables for at least 8.5 years.
@@ 13.1 General payload description
The PLATO payload has two major science drivers constraining its design. Compared to previous missions like Kepler, PLATO has been optimised to observe stars bright enough to allow radial velocity measurements on ground to complement the photometric measurements in space. Since such bright stars are relatively scarce on the sky, a very large total field of view of 2132 deg2 was required. Furthermore, to be able to observe in a given target field the brightest as well as fainter stars, the instrument photon collecting area is split into 26 individual cameras. The 24 cameras which are used exclusively for science operation are termed "normal" cameras whereas the two additional cameras, used also for fine-pointing on top of science operation, are termed "fast" cameras. With this configuration PLATO will be able to obtain very stable, high precision, long-time photometric data of a large number of bright stars. Combining the information provided by the planetary transits, with the stellar data derived by asteroseismology techniques and the radial velocity measurements obtained from the ground, PLATO will provide a high-precision characterisation of the radius, mass and age of exoplanetary systems.
@@
Table 8 summarises the main data of the instrument. Each 20 cm class camera (12 cm entrance pupil) has a total field of view (FoV) of 1037 square degrees [see 271]. The 24 normal cameras are organised on the optical bench of the spacecraft in 4 groups of 6 co-aligned cameras each. The sky fields of the four camera groups are offset with respect to a central line of sight by 9.2 degrees, pointing towards the four corners of a square of 13.0 deg, in order to widen the total overlapping FoV of the instrument. With this approach the number of brightest stars within SNR requirements is augmented with respect to a full coalignment of the whole set of cameras. Figure 25 illustrates how many cameras observe the same part of the target field for a pointing. The total instrument field therefore is about 2132 deg² (with rounded edges). Figure 26 shows the Structural and Thermal Model (STM) of the payload during integration at OHB, ready for the qualification activities of the Payload Module.
@@
| Total FoV | 2132 deg² |
| FoV by 24 CAMs | 294 deg² |
| FoV by 18 CAMs | 171 deg² |
| FoV by 12 CAMs | 796 deg² |
| FoV by 6 CAMs | 871 deg² |
@@
| PDE | 3 arcsec HCA over 3 months in the directional direction |
| 6 arcsec HCA over 3 months in the rotational direction | |
| PRE | 3 arcsec HCA in the directional direction |
| 6 arcsec HCA in the rotational direction | |
| RPE | 1 arcsec HCA over 25s in the directional direction |
| 2 arcsec HCA over 25s in the rotational direction | |
@@
| Peak power consumption | 890 W |
@@
| Normal electronic chain | 12 DPUs: 1 for 2 normal cameras |
| 2 normal AEUs: 1 for 12 normal cameras | |
@@
PDE: Attitude Pointing Drift Error; PRE: Attitude Pointing Repeatability Error; RPE: Relative Attitude Pointing Error.
FoV: Total Field of View; HCA: half cone angle; DPU: Data Processing Unit; AEU: Ancillary Electronic Unit.